# references.bib

### v1

@@ Abstract
Spiking Neural Networks (SNNs) are biologically-inspired deep neural networks that efficiently extract temporal information while offering promising gains in terms of energy efficiency and latency when deployed on neuromorphic devices. However, SNN model parameters are sensitive to temporal resolution, leading to significant performance drops when the temporal resolution of target data at the edge is not the same with that of the pre-deployment source data used for training, especially when fine-tuning is not possible at the edge. To address this challenge, we propose three novel domain adaptation methods for adapting neuron parameters to account for the change in time resolution without re-training on target time-resolution. The proposed methods are based on a mapping between neuron dynamics in SNNs and State Space Models (SSMs); and are applicable to general neuron models. We evaluate the proposed methods under spatio-temporal data tasks, namely the audio keyword spotting datasets SHD and MSWC as well as the image classification NMINST dataset. Our methods provide an alternative to - and in majority of the cases significantly outperform - the existing reference method that simply scales the time constant. Moreover, our results show that high accuracy on high temporal resolution data can be obtained by time efficient training on lower temporal resolution data and model adaptation.
Index Terms:
domain adaptation, temporal resolution, spiking neural networks (SNNs), state-space models (SSMs).
I Introduction
@@
In this paper, we are interested in the scenario where the temporal resolution of the source data that the model has been trained on and the target data that the model is to be used on is different. We investigate both the scenario of Fine-to-Coarse deployment, where the change is from high to low resolution, and Coarse-to-Fine deployment, where the change is from low to high resolution. We consider the setting where it is not possible to fine-tune the model at the edge on the data with the target temporal resolution, for instance when real-time inference is needed on the current sample from target domain. Hence, we consider temporal resolution domain adaptation with no re-training on the target dataset, see Figure 1.
@@
A possible application of Coarse-to-Fine deployment is encountered in communication constrained distributed scenarios with a central learning unit. In such scenarios, sensor data is sent to a central learning unit with reduced temporal resolution due to bandwidth or power limitations [chen2021distributed, park2021communication]. As a result, the central model is trained on coarse data and then deployed to edge devices, where higher-resolution data is available for inference. A possible application of Fine-to-Coarse deployment is encountered when the sensors have to switch to low sampling rates to reduce power consumption [Dieter_2005, Ankur2004]. This leads to having target data at edge with coarser temporal resolution than the source data used for training the deployed models at the central learning unit.
We consider the above temporal domain adaptation problems under Spiking Neural Networks (SNN) models. SNNs are biologically-inspired neural networks with stateful recurrent neurons and binary spike-based communication. SNNs provide an attractive framework for processing data with encoded temporal information due to their ability to represent and learn spatiotemporal features [bittar2022surrogate, yik2024neurobench, Braille_SNN, eshraghian2023trainingspikingneuralnetworks]. Like most machine learning algorithms, the current mainstream pipeline for SNNs relies on training on high-power computational units, mostly GPUs, and then deploying and running on low-power edge devices [10177729]. SNNs, when deployed at neuromorphic edge hardware, are especially valuable because they are designed to mimic the brain’s energy-efficient processes, offering significant reductions in power consumption and latency [9395703, caccavella2023lowpower, Braille_SNN]. SNNs are particularly effective in event-based systems, where event-based sensors, e.g. event-cameras record changes only as they occur, rather than at fixed time intervals, leading to optimized energy use and reduced data redundancy [Gallego2022].
@@
  • The results show that our proposed methods provide significant performance gains compared to using no temporal resolution domain adaptation method or using the existing benchmark method from the literature.
@@
II Related Work
The Leaky-Integrate-and-Fire (LIF) neuron model is one of the most simple and widely used neuron models in the context of SNNs. When input stimuli in form of spikes are applied to the neuron, the membrane potential increases and when a certain threshold is reached it produces an output spike. On the other hand in absence of input stimuli the membrane potential decays exponentially [bittar2022surrogate], see Figure 2. The vanilla LIF neuron has only one learnable parameter, which is the membrane potential decay (the leak), and this parameter has explicit exponential dependence on the time step length of the data. Hence when dealing with change of temporal resolution, intuitive way to adapt the LIF neuron dynamic is to scale the leak parameter according to the change. In [he2020comparing] the authors explore this method of temporal domain adaptation using the simple LIF neuron. On the other hand, in [caccavella2023lowpower], the authors use Integrate-and-Fire (IF) neuron which is a special simpler case of LIF neuron without leak. Hence alternative to parameters adaptation, they employ a graded spike network and different normalization techniques for domain adaptation with promising results. However, in order to capture more complex temporal information in data, more complex neuron models have been used [bittar2022surrogate][izhikevich2003simple], in which parameter might not have explicit dependence on the time-length of the data, or even any dependence at all. In this case, it is unclear how these parameters need to be scaled to depict the change of temporal resolution, and to the best of our knowledge, this issue has not been thoroughly addressed in the existing literature.
Figure 2: Feedforward SNN model architecture consisting of four layers, each with arbitrary number of neurons. An example LIF neuron dynamics given as a zoom-in on a neuron.
@@ II Related Work
Interesting intersection of work is emerging between SNNs and deep SSMs. One common line of work is embedding spiking dynamics in deep SSMs where at each time step the output activity of the SSM is processed by a firing threshold function [stan2023learning] [bal2024rethinkingspikingneuralnetworks]. Alternatively, in [du2024spiking] they combine SNN with SSMs by adding a spiking layer to the SSM. Nevertheless, the challange of drawing lines of correspondence between SNNs and linear SSMs is still present due to the non-linearity when it comes to SNNs. Our proposed methods for temporal domain adaptation for SNNs are based on this drawn parallel and the theory developed for adaptation in SSMs.
III Problem Statement
In Section III-A we present popular neuron models commonly used in the SNN literature to establishes a foundation for the generalized neuron model in Section III-B. We outline the temporal domain adaptation problem this article focuses on in Section III-C.
III-A Popular spiking neuron models
The discrete-time LIF neuron is defined as [bittar2022surrogate]
$\!u[t+1]$ $=\alpha(u[t]-\theta s_{out}[t])+(1-\alpha)(\bm{W}\bm{s}_{in}[t]+\bm{V}\bm{s}_{out}[t])$ (1a)
$s_{out}[t]$ $=g_{\theta}(u[t])=\begin{cases}1\text{ if }u[t]\geq\theta\\ 0\text{ otherwise}\end{cases}$ (1b)
where $\bm{s}_{in}\in\mathbb{Z}^{L_{b}\times 1}$ are the input spikes from the neurons in the previous layer and $\bm{W}\in\mathbb{R}^{1\times L_{b}}$ is the weight matrix associated with their connections to the neuron of observation, while $\bm{s}_{out}\in\mathbb{Z}^{L_{c}\times 1}$ are the output spikes of all the other neurons in the current layer and $\bm{V}\in\mathbb{R}^{1\times L_{c}}$ with zero diagonal is their connections to the neuron of observation. Here $L_{c}$ and $L_{b}$ are the number of neurons in the current and the previous layer, respectively. Furthermore, $g_{\theta}$ describes the spiking behaviour with firing threshold $\theta\in\mathbb{R}^{1\times 1}$; $\alpha=\exp(\Delta/\tau_{mem})$ represents the decay of the membrane potential over time such that $\Delta$ corresponds to the duration of a timestep; and $\tau_{mem}$ is a membrane time constant. Usually, $\alpha$ is a learnable parameter initialized uniformly and trained within a given range.
The adaptive LIF (adLIF) neuron, an extension of the LIF, introduces a recovery variable $v[n]$ of the membrane potential which results in exhibiting an input adaptive behavior [bittar2022surrogate]. In discrete-time, it is defined as follows:
@@ III-A Popular spiking neuron models
where the adjustable parameters are $\alpha$, as defined before, $\beta=\exp(\Delta/\tau_{syn})$ with adaptation constant $\tau_{syn}$, and $a$ and $b$ both without explicit dependence on the time step length $\Delta$. All parameters here are generally learnable during training within given ranges. Note, in both LIF and adLIF neurons, the weight vector $\bm{V}$ corresponds to recurrent connections between the neurons in the same layer. Recurrence represented by $\bm{V}$ is a characteristic of the network rather than the neuron.
III-B General spiking neuron model
In this section we introduce a discrete-time general n-dimensional spiking neuron model which generalizes the one-dimensional LIF neuron (1), and the two-dimensional adLIF neuron (2). This generalized model is central to the proposed methods of this paper. It is defined as
@@ III-B General spiking neuron model
We note that the LIF neuron (1) with one state variable $\bm{v}=u$ is a special case of the general neuron model where the matrices are in fact scalars with $\bm{H}_{v}=\alpha,\bm{H}_{f}=-\alpha\theta,\bm{H}_{i}=1-\alpha,$ and $\bm{H}_{r}=1-\alpha.$
Similarly, we note that the adLIF neuron can be expressed as
@@
III-C Problem Formulation - Temporal resolution domain adaptation for general spiking neuron model
III-C1 Event-based data
@@ III-C3 Model adaptation
IV Methods
As a background for our work, in Section IV-A, we begin by expressing the SNN neuron model as a non-linear SSM model. This provides a foundation for Section IV-B, where we propose an approximate correspondence between the SNN neuron model and linear SSM model. The proposed temporal resolution domain adaptation methods, presented in Section IV-C, are based on this approximate correspondence. In Section IV-D we present a simple benchmark method.
IV-A Preliminaries- Correspondence between SNNs and non-linear SSMs
@@
IV-B Correspondence between SNNs and linear SSMs
@@
IV-C Proposed Methods for Temporal Resolution Domain Adaptation
In this section, we define our proposed temporal resolution domain adaptation methods for the generalized SNN neuron model in (3). As explained in Section III-C, the proposed methods suggest a mapping ${\cal{M}}(.)$ from the source domain with temporal resolution $\Delta_{S}$ to the target domain with temporal resolution $\Delta_{T}$, where $\rho=\frac{\Delta_{S}}{\Delta_{T}}$.
@@ IV-C Proposed Methods for Temporal Resolution Domain Adaptation
In the Integral and Expectation adaptation methods, $\left(\bm{H}_{v,(S)}\right)^{\rho}$ represents a matrix raised to a possibly fractional power which can be computed using the Schur–Padé algorithm [MatrixFractionalPower]. By restricting the parameters of the adLIF neuron (2) to the ranges $\alpha,\beta,a\in(0,1)$ and $b\in(0,2)$, the possibly fractional power $\rho=1/\bar{\rho}$ of the real matrix $\bm{H}_{v,(S)}$ used in our experiments remains real-valued [MatrixFractionalPower].
IV-D A Simple Benchmark Method: Time-constant Adaptation Method
@@
To illustrate the method, we now look into the popular adaptive LIF neuron from Section III-A. In the adaptive LIF neuron, all of the three types of elements are present. Here, $\alpha_{(T)}=\alpha_{(S)}^{\rho}$ and $\beta_{(T)}=\beta_{(S)}^{\rho}$ due to their explicit exponential dependence on the step length $T$. The other parameters are either scaled accordingly due to the dependence with the already scaled parameters, for example $-(1-\alpha_{(T)})$ or kept the same (in the case of the parameters $a$ and $b$).

### v2

# Zero-Shot Temporal Resolution Domain Adaptation for Spiking Neural Networks

@@ Abstract
Spiking Neural Networks (SNNs) are biologically-inspired deep neural networks that efficiently extract temporal information while offering promising gains in terms of energy efficiency and latency when deployed on neuromorphic devices. However, SNN model parameters are sensitive to temporal resolution, leading to significant performance drops when the temporal resolution of target data during deployment is not the same with that of the pre-deployment source data used for training, especially when fine-tuning with new data is not possible during deployment. To address this challenge, we propose three novel domain adaptation methods for adapting neuron parameters to account for the change in time resolution without re-training on target time-resolution. The proposed methods are based on a mapping between neuron dynamics in SNNs and State Space Models (SSMs) and are applicable to general neuron models. We evaluate the proposed methods under spatio-temporal data tasks, namely the audio keyword spotting datasets SHD and MSWC as well as the neuromorphic image classification NMINST dataset. Our methods provide an alternative to - and in majority of the cases significantly outperform - the existing reference method that simply scales the time constant. Moreover, our results show that high accuracy on high temporal resolution data can be obtained by time efficient training on lower temporal resolution data and model adaptation.
keywords:
domain adaptation, temporal resolution, neuromorphic, spiking neural networks (SNNs), state-space models (SSMs).
journal: Neural Networks
[…]
1 Introduction
@@
In this paper, we are interested in the scenario where the temporal resolution of the source data which the model has been trained on, and the target data which the model is to be used on is different. We investigate both the scenario of Fine-to-Coarse deployment, where the change is from high to low resolution, and Coarse-to-Fine deployment, where the change is from low to high resolution. We consider the setting where it is not possible to fine-tune the model at the edge on the data with the target temporal resolution, for instance when real-time inference is needed on the current sample from target domain. Hence, we consider temporal resolution domain adaptation with no re-training on the target dataset, see Figure 1.
@@
Applications of these scenarios appear at different edge processing scenarios, where small resource-constrained devices with different sensing and computing capabilities must process their data at the edge of a wireless system with varying degrees of support from a central processing or coordination unit [chen2021distributed, park2021communication]. A possible application of Coarse-to-Fine deployment is encountered in communication constrained distributed scenarios with a central learning unit. In such scenarios, sensor data is sent to a central learning unit with reduced temporal resolution due to bandwidth or power limitations [chen2021distributed, park2021communication]. As a result, the central model is trained on coarse data and then deployed to edge devices, where higher-resolution data is available for inference. A possible application of Fine-to-Coarse deployment is encountered when the sensors have to switch to low sampling rates to reduce power consumption [Dieter_2005, Ankur2004]. This leads to having target data at edge with coarser temporal resolution than the source data used for training the deployed models at the central learning unit.
We consider the above temporal domain adaptation problems under Spiking Neural Networks (SNN) models. SNNs are biologically-inspired neural networks with stateful recurrent neurons and binary spike-based communication [MAASS19971659]. SNNs provide an attractive framework for processing data with encoded temporal information due to their ability to represent and learn spatiotemporal features [bittar2022surrogate, yik2024neurobench, Braille_SNN, eshraghian2023trainingspikingneuralnetworks, ma2025spikingneuralnetworkstemporal]. Like most machine learning algorithms, the current mainstream pipeline for SNNs relies on training on high-power computational units, mostly GPUs, and then deploying and running on low-power edge devices [10177729]. SNNs, when deployed at neuromorphic edge hardware, are especially valuable because they are designed to mimic the brain’s energy-efficient processes, offering significant reductions in power consumption and latency [9395703, caccavella2023lowpower, Braille_SNN]. SNNs are particularly effective in event-based systems, where event-based sensors, e.g. event-cameras record changes only as they occur, rather than at fixed time intervals, leading to optimized energy use and reduced data redundancy [Gallego2022].
@@
  • The results show that in a wide range of scenarios, e.g. coarse-to-fine scenarios for all the above datasets, our proposed methods provide significant performance gains compared to using no temporal resolution domain adaptation method or using the existing benchmark method from the literature.
@@
2 Related Work
The Leaky-Integrate-and-Fire (LIF) neuron model is one of the most simple and widely used neuron models in the context of SNNs. When input stimuli in form of spikes are applied to the neuron, the membrane potential increases and when a certain threshold is reached it produces an output spike. On the other hand, in the absence of input stimuli, the membrane potential decays exponentially [bittar2022surrogate], see Figure 2. The vanilla LIF neuron has only one learnable parameter, which is the membrane potential decay (the leak), and this parameter has explicit exponential dependence on the time step length of the data. Hence when dealing with change of temporal resolution, intuitive way to adapt the LIF neuron dynamic is to scale the leak parameter according to the change. In [he2020comparing] the authors explore this method of temporal domain adaptation using the simple LIF neuron. On the other hand, in [caccavella2023lowpower], the authors use Integrate-and-Fire (IF) neuron which is a special simpler case of LIF neuron without leak. Hence alternative to parameters adaptation, they employ a graded spike network and different normalization techniques for domain adaptation with promising results. However, in order to capture more complex temporal information in data, more complex neuron models such as the adaptive LIF (adLIF) neurons have shown superior performance [bittar2022surrogate, baronig2025advancingspatiotemporalprocessingspiking, adLIFNature, ma2025spikingneuralnetworkstemporal]. In the case of more complex neuron models however, the parameter might not have explicit dependence on the time-length of the data, or even any dependence at all.
In this case, it is unclear how these parameters need to be scaled to depict the change of temporal resolution, and to the best of our knowledge, this issue has not been thoroughly addressed in the existing literature.
Figure 2: Feedforward SNN model architecture consisting of four layers, each with arbitrary number of neurons. An example neuron dynamics given as a zoom-in on a neuron.
@@ 2 Related Work
Interesting intersection of work is emerging between SNNs and deep SSMs. One common line of work is embedding spiking dynamics in deep SSMs where at each time step the output activity of the SSM is processed by a firing threshold function [stan2023learning, bal2024rethinkingspikingneuralnetworks, karilanova2025statespacemodelinspiredmultipleinput]. Alternatively, in [du2024spiking] they combine SNN with SSMs by adding a spiking layer to the SSM. Nevertheless, the challange of drawing lines of correspondence between SNNs and linear SSMs is still present due to the non-linearity when it comes to SNNs. Our proposed methods for temporal domain adaptation for SNNs are based on this drawn parallel and the theory developed for adaptation in SSMs.
3 Problem Statement
In Section 3.1 we present a popular neuron model commonly used in the SNN literature to establish a foundation for the generalized neuron model in Section 3.2. We outline the temporal domain adaptation problem that this article focuses on in Section 3.3.
3.1 Popular spiking neuron model
The discrete-time adaptive Leaky-Integrate and Fire (adLIF) neuron is defined as [bittar2022surrogate]
@@ 3.1 Popular spiking neuron model
where $\bm{s}_{in}\in\mathbb{Z}^{L_{b}\times 1}$ are the input spikes from the neurons in the previous layer and $\bm{W}\in\mathbb{R}^{1\times L_{b}}$ is the weight matrix associated with their connections to the neuron of observation, while $\bm{s}_{out}\in\mathbb{Z}^{L_{c}\times 1}$ are the output spikes of all the other neurons in the current layer and $\bm{V}\in\mathbb{R}^{1\times L_{c}}$ with zero diagonal is their connections to the neuron of observation. Here, $L_{c}$ and $L_{b}$ are the number of neurons in the current and the previous layer, respectively. Furthermore, $g_{\theta}$ describes the spiking behavior with firing threshold $\theta\in\mathbb{R}^{1\times 1}$; $\alpha=\exp(\Delta/\tau_{mem})$ represents the decay of the membrane potential over time such that $\Delta$ corresponds to the duration of a timestep and $\tau_{mem}$ is a membrane time constant; $\beta=\exp(\Delta/\tau_{syn})$ represents the decay of the adaptive variable with adaptation constant $\tau_{syn}$; the parameters $a$ and $b$ are without explicit dependence on the discrete time step length $\Delta$. All parameters $\alpha,\beta,a,b$ are generally learnable during training within given ranges.
Note, the weight vector $\bm{V}$ corresponds to recurrent connections between the neurons in the same layer. Recurrence represented by $\bm{V}$ is a characteristic of the network rather than the neuron.
3.2 General spiking neuron model
In this section we introduce a discrete-time general n-dimensional spiking neuron model which generalizes the two-dimensional adLIF neuron (1). This generalized model is central to the proposed methods of this paper. It is defined as
@@ 3.2 General spiking neuron model
We note that the adLIF neuron can be expressed as
@@
3.3 Problem Formulation - Temporal resolution domain adaptation for general spiking neuron model
3.3.1 Event-based data
@@ 3.3.3 Model adaptation
4 Methods
As a background for our work, in Section 4.1, we begin by expressing the SNN neuron model as a non-linear SSM model. This provides a foundation for Section 4.2, where we propose an approximate correspondence between the SNN neuron model and linear SSM model. The proposed temporal resolution domain adaptation methods, presented in Section 4.3, are based on this approximate correspondence. In Section 4.4 we present a simple benchmark method.
4.1 Preliminaries - Correspondence between SNNs and non-linear SSMs
@@
4.2 Preliminaries - Correspondence between SNNs and linear SSMs
@@
4.3 Proposed Methods for Temporal Resolution Domain Adaptation
In this section, we define our proposed temporal resolution domain adaptation methods for the generalized SNN neuron model in (2). As explained in Section 3.3, the proposed methods suggest a mapping ${\cal{M}}(.)$ from the source domain with temporal resolution $\Delta_{S}$ to the target domain with temporal resolution $\Delta_{T}$, where $\rho=\frac{\Delta_{T}}{\Delta_{S}}$.
@@ 4.3 Proposed Methods for Temporal Resolution Domain Adaptation
In the Integral and Expectation adaptation methods, $\left(\bm{H}_{v,(S)}\right)^{\rho}$ represents a matrix raised to a possibly fractional power which can be computed using the Schur–Padé algorithm [MatrixFractionalPower]. In general, fractional powers of a real-valued matrix may be complex-valued. By restricting the parameters of the adLIF neuron (1) to the ranges $\alpha,\beta,a\in(0,1)$, the possibly fractional power $\rho=1/\bar{\rho}$ of the real matrix $\bm{H}_{v,(S)}$ used in our experiments remains real-valued [MatrixFractionalPower].
4.4 A Simple Benchmark Method: Time-constant Adaptation Method
@@
To illustrate the method, we now look into the popular adaptive LIF neuron from Section 3.1. In the adaptive LIF neuron, all of the three types of elements are present. Here, $\alpha_{(T)}=\alpha_{(S)}^{\rho}$ and $\beta_{(T)}=\beta_{(S)}^{\rho}$ due to their explicit exponential dependence on the step length $T$. The other parameters are either scaled accordingly due to the dependence with the already scaled parameters, for example $-(1-\alpha_{(T)})$ or kept the same (in the case of the parameters $a$ and $b$).